OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = 6x2 – 24x + 32? Answer a.(–2, 8) b.(2, 8) c.(2, –8) d.(–2, –8)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

y=6(x^2-4x)+32 y=6(x-2)^2-24+32 y=6(x-2)^2+8 so, the vertex is (2,8) is B!!!!!